14x+6y=24
6x+14y=12

Answer by mananth(16949)

14x+6y=24.............1
6x+14y=12.............2
add the two equations
we get 20x+20y=36
/4
5x+5y=9
subtracrt (2) from (1)
8x-8y=12
/4
2x-2y=3
5 x + 5 y = 9 .............1
2 x -2 y = 3 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 5
10 x 10 y = 18
10 x -10 y = 15
Add the two equations
20 x = 33
/ 20
x = 33/20
plug value of x in (1)
5 x + 5 y = 9
33/4 + 5 y = 9
5 y = 9 -33/4
5 y = 3/4
y = 3/20